OpenStudy (anonymous):

how do i divide 6 radical 3 over 2?

OpenStudy (kropot72):

Is this right?\[(6\times \sqrt{3})/2\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[{{6\sqrt{3}}\over2}={{3\sqrt{3}}}\]

OpenStudy (anonymous):

thanks!